# Effect Identification and Unit Categorization in the Multi-Score Regression Discontinuity Design with Application to LED Manufacturing

### v1

@@ Abstract
The RDD (regression discontinuity design) is a widely used framework for identification and estimation of causal effects at a cutoff of a single running variable. Practical settings, in particular those encountered in production systems, often involve decision-making defined by multiple thresholds and criteria. Common MRD (multi-score RDD) approaches transform these to a one-dimensional design, to employ identification and estimation results. However, this practice can introduce non-compliant behavior. We develop theoretical tools to identify and reduce some of this “fuzziness” when estimating the cutoff-effect on compliers of sub-rules. We provide a sound definition and categorization of unit behavior types for multi-dimensional cutoff-rules, extending existing categorizations. We identify conditions for the existence and identification of the cutoff-effect on complier in multiple dimensions, and specify when identification remains stable after excluding nevertaker and alwaystaker. Further, we investigate how decomposing cutoff-rules into simpler parts alters the unit behavior. This allows identification and removal of non-compliant units potentially improving estimates. We validate our framework on simulated and real-world data from opto-electronic semiconductor manufacturing. Our empirical results demonstrate the usability for refining production policies. Particularly we show that our approach decreases the estimation variance, highlighting the practical value of the MRD framework in manufacturing.
@@ 1 Introduction
In Economics and Management – in particular Marketing, Finance, Social sciences, and many other subfields – decision-making often hinges on multiple criteria. Examples include targeted marketing, education policies, political advertising.
The complexity of multiple criteria has led to the development of formal frameworks such as Multi-Criteria Decision Analysis, which structure these trade-offs and provide methods for ranking alternatives. Yet, as decisions become embedded in policy environments, a new challenge emerges: understanding the actual impact of those decisions. Here, the focus shifts from choosing among options to rigorously evaluating outcomes. This is where causal inference enters, and with it, designs like Regression Discontinuity Design (RDD).
@@
RDD has emerged as a powerful quasi-experimental tool for identifying causal effects across diverse fields including economics (Hartmann et al., 2011; Card et al., 2015; Flammer, 2015; Calvo et al., 2019), public policy (Lee, 2008) and the social sciences (Angrist and Lavy, 1999), whenever a policy is in place that leads to a jump in treatment probability at a certain cutoff. Its appeal lies in its ability to deliver credible causal estimates under minimal assumptions. Particularly, RDD can be applied even if typical causal inference assumptions such as unconfoundedness Rubin (1974) or positivity (Austin, 2011) do not hold.
@@
The classic RDD utilizes discontinuities in a single treatment assignment rule that is based on an observed running or score variable. When correctly specified, this setting enables identification of average treatment effects for units near the cutoff. Thus, under mild conditions, it enables to counter-factually reason about the decision policy in place in a small neighborhood around the boundary (Cattaneo et al., 2019; Imbens and Lemieux, 2008; Lee and Lemieux, 2010).
@@
As motivated above, in many real-world contexts – particularly in the scope of industrial, operational, or engineering systems – treatment decisions are typically based on multiple criteria (Sabaei et al., 2015). The logical extension of the classical RDD to multi-score RDD (MRD), in which treatment decisions are taken under a joint logic of multiple score variables satisfying a threshold condition has been proposed by Papay et al. (2011). Influential surveys that propose estimators and empirically study them, include Reardon and Robinson (2012), Wong et al. (2013) and Porter et al. (2017). Particularly in test score-based (An et al., 2024) and geographical (Keele and Titiunik, 2015) applications MRD has been successfully applied in many empirical studies. However, identification strategies of MRD effects for multiple cutoffs in an arbitrarily connected logical rule remains unclear in the literature.
@@
This paper addresses the identification and estimation challenges that arise in the context of multi-score RDD. We develop new identification results tailored for multi-score RDD settings. Specifically, we provide a formal framework that defines and categorizes behavioral types (“complier”, “alwaystaker”, “nevertaker”, “defier”) of units and addresses the inheritance of these types when changing the treatment assignment. We investigate general boolean-type cutoff rules, such as "AND"-type and "OR"-type rules, commonly observed in MRD environments and analyze how their properties influence local identification. Further, we show how the complier effect can be efficiently identified using these categories. We then showcase our results on a dataset of a light-emitting diode (LED) production facility, where each batch undergoes evaluation based on several metrics – such as mean color point measurements or share of units per lot that fulfill a certain quality specification.
@@
Our application of the proposed strategies to data from the LED manufacturing industry demonstrates the utility of our framework in a real-world problem, where treatment assignment is based on multiple quality indicators. We combine our framework with recently proposed estimators that use ML-based adjustments to further reduce variance in estimation (Noack et al., 2024). The estimation of the causal effect of a threshold-based production guides the optimal tuning of the decision threshold to improve manufacturing outcomes. Specifically, managers need to ascertain how marginal adjustments to these thresholds affect outcomes like yield rates and thus the overall production efficiency. We complement our empirical results with a simulated study using a semi-synthetic clone of the production environment, highlighting the value of multi-score RDD for counterfactual analysis and manufacturing policy design.
@@
This study contributes to the operations research literature by bridging causal machine learning advancements with practical decision-making in manufacturing systems. Our theoretical results on effect identification in MRD open new paths for its application in decision making contexts. By accommodating complex, multi-criteria assignment mechanisms, our framework aligns with the realities of operational decision-making, offering tools for more informed and effective policy design.
@@
Our paper is organized as follows. Section 2 reviews related literature on RDD and its extensions to multi-score settings. Section 3 presents our theoretical framework for identification under general Boolean assignment rules. Section 4 details our empirical application in the scope of opto-electronic semiconductor manufacturing. Section 5 provides empirical results of our proposed identification strategies in both synthetic and semi-synthetic environments. Section 6 discusses results from a real data application, highlighting the implications for threshold calibration and production efficiency. Section 7 concludes the paper.
@@ 2 Related Literature
The Regression Discontinuity Design dates back to a study on scholarship programs by Thistlethwaite and Campbell (1960). In the following decades, multiple studies employing this quasi-experimental approach gained popularity in the social sciences, such as Angrist and Lavy (1999) and (Black, 1999). Building on this pioneering empirical work, Hahn et al. (2001) first formalized the RD design within the potential-outcomes framework and delivered non-parametric identification and estimation results for both sharp and fuzzy designs, firmly establishing RD as a core tool of modern econometrics.
@@
Impactful surveys including Imbens and Lemieux (2008), Lee and Lemieux (2010) and Cattaneo et al. (2019), who synthesize clear guidance on estimation choices for the local linear regression that serves as estimator, such as bandwidth selection, polynomial order, and diagnostic checks. There were multiple key theoretical contributions recently. Calonico et al. (2014) developed bias-corrected point estimates and robust confidence intervals that improve coverage accuracy and have become the empirical default for RD inference. Calonico et al. (2019b) study the addition of covariates to the local regression in order to reduce the estimator’s variance. Calonico et al. (2019a) study the optimal bandwidth choice for robust RDD estimators. Further extensions that include ML-based covariate adjustments in order to reduce variance in estimation have brought RDD into the toolbox of causal ML researchers (Kreiss and Rothe, 2022; Arai et al., 2025; Noack et al., 2024).
@@
The idea of extending RDD to a multi-score setting dates back to a study from Papay et al. (2011) showing that causal effects can be identified along a two-dimensional cutoff frontier when program eligibility depends jointly on multiple scores. Reardon and Robinson (2012) first systematically collected a set of MRD identification strategies when the running variables jointly cross their cut-offs as a multidimensional “frontier”, translating to an “AND”-connection of rules. They develop both frontier-projection and distance-based estimators, and highlight the steep data-density and bandwidth challenges that arise as the dimensionality of the assignment space grows. Stepping from conceptual proposals to systematic evaluation, Wong et al. (2013) performed extensive simulation studies and demonstrated that estimators which weight observations along the full cutoff frontier achieve lower bias and higher efficiency than naive one-dimension-at-a-time approaches, highlighting the need to exploit the joint assignment rule in practice. In another extensive study, Porter et al. (2017) derive similar findings and give practical advice for MRD settings. An important special case, the geographic RD, tailored to evaluate policies along irregular (state) borders was introduced by Keele and Titiunik (2015). Recently, also more theoretically driven work appeared. Choi and Lee (2018) allow for partial treatment effects when only some running variables cross their cut-offs and introduce thin-plate-spline estimators that flexibly recover these heterogeneous impacts along a high-dimensional frontier. Choi and Lee (2023), building upon earlier work by Imbens and Zajonc (2009), delivered insights on two-dimensional MRD with limited compliance and thus first deriving estimation implications for complier effects in fuzzy MRD.
@@
Recent work has also introduced more advanced MRD estimators based on a minimax approach (Imbens and Wager, 2019), on decision-tree methods (Liu and Qi, 2024), and on fully multivariate local-polynomial estimators (Sawada et al., 2025). Finally, Cattaneo et al. (2025) discuss the statistical properties of important MRD estimators in more detail, giving practical recommendations.
@@
Recent work has proposed the application of the regression discontinuity design in business decision making (Hünermund et al., 2021) or in production operations management (Ho et al., 2017; Mithas et al., 2022). Calvo et al. (2019) use a one-dimensional fuzzy RD design to estimate the effect of procurement officers on budget overruns and delays in public infrastructure projects. Further applications include Leung et al. (2020) who study the effects of the labor unionization of firms on the operating performance of their suppliers. RDD has also been employed to assess the impact of regulation changes on compliance in the car industry (Hu et al., 2021). However, to the best of our knowledge, there are no direct applications of RDD nor MRD to production policies.
@@
This paper contributes to the literature in two main ways: Firstly, we derive theoretical results yielding new results on unit categorization and the identification in MRD that are not limited to, but particularly useful, when dealing with MRD in complex, multi-dimensional decision making. Secondly, we apply our theoretical results in a real-data application, providing a framework that supports the applicability of MRD in operations policy making by deriving implications on a real-world decision policy.
@@ 3 General Identification Strategies in Multi-score RDD
In this section, we first introduce the basics of the (Multi-Score) RDD design. Having established this, we give a formal definition of common behavior types of units (e.g., complier, defier) in multi-score, two-stage decision settings that employ cutoff-rules for the initial treatment assignment. We derive an identification result using these unit categories. In particular, we show that under certain assumptions, the identification does not depend on subsets of unit types with constant response. Additionally, we derive rules for the inheritance of unit types as the assignment rules get more complex.
@@ 3.1 Regression Discontinuity Design
Recent sources rely on the groundwork of Hahn et al. (2001), which provides a formal justification for effect identification and estimation using RDD. RDD consists of three key ingredients: A score $X$ that defines a certain characteristic of the units, a cutoff $c$ that splits the support of the score into two groups and a treatment $D$, which is assigned to one of the groups based on their score and the cutoff, $D_{i}=\operatorname{\mathbbm{1}}[X_{i}\geq c]$ (Cattaneo et al., 2019).
Typically, a researcher is interested on the average treatment effect at the cutoff $c$, namely $\mathbb{E}[Y(1)-Y(0)\,|\,X=c]$ with $Y(1)$ and $Y(0)$ being the potential outcomes of individuals according to Rubin (2005). The identification of the effect relies on the following key assumption:
@@
The conditional mean of the potential outcomes $\mathbb{E}[Y_{i}(d)\mid X_{i}=x]$ for $d\in\{0,1\}$ is continuous at the cutoff level $c$.
This assumption implies that units that receive very similar score on opposite sides of the cutoff are comparable to each other in all relevant aspects (Hahn et al., 2001). To ensure that there is no selection on gains, we have to assure that units around the threshold are not able to perfectly manipulate their score, formally, $\mathbb{E}[Y(1)-Y(0)\mid D,\,X=x]=\mathbb{E}[Y(1)-Y(0)\mid X=x]$ has to hold for $x\in(c-\varepsilon,c+\varepsilon)$ and small $\varepsilon>0$ (Chernozhukov et al., 2024).
Under these relatively mild assumptions, RDD provides inference around the threshold as credible as in a randomized experiment Lee (2008). The average treatment effect at the threshold
@@ 3.2 Problem description
While standard RDD is suitable for analyzing effects of a single score variable at a cutoff, real-world treatment decisions often involve multiple score variables. For example, in the case when treatment $T_{i}=1$ is assigned to the unit $i$ only if both criteria $X_{1,i}>c_{1}$ and $X_{2,i}>c_{2}$ are met. Oftentimes, such multi-score cases are reduced to a standard RDD setting by transforming the involved variables into a single score. The identification and estimation of the MRD effect can then be justified by applying theoretical results from the one-dimensional case. One drawback, however, is the reduced interpretability of the transformed score. Thus, the question arises under which precise conditions a treatment effect is existent and identifiable in a multi-score setting. In particular, a well-defined multidimensional cutoff effect requires a certain independence of the directions from which the cutoff is approached.
@@
In applications, one is usually interested in optimizing the treatment assignment $T_{i}$ to improve the outcome $Y_{i}$. Critical systems might only allow for gradual optimization of $T_{i}$. In such cases, a part of the assignment rule $T_{i}$, e.g., $H_{i}\coloneqq\operatorname{\mathbbm{1}}[X_{2,i}>c_{2}]$, is accepted as given, while the remaining part, e.g., $G_{i}\coloneqq\operatorname{\mathbbm{1}}[X_{1,i}>c_{1}]$, is subjected to optimization. Thus, the effect of a subrule $G$ of $T$ at its cutoff can provide insights on how to improve $T$. In setups where the treatment assignment $T$ is seen as a mere recommendation, the actual treatment $D$ is allowed to deviate from it. In RDD, this case is usually referred to as “fuzzy”. All in all, this results in a hierarchy $G\to T\to D$ of decision rules. Each transition potentially obfuscates the assessment of $G$ – in our case, the assessment of $c_{1}$. The reason is clear: There might be units that do not comply with $G$ at each level of the hierarchy. On the other hand, knowledge of the compliance behavior of a unit $i$ with respect to the rules (G,D)(G,\,D) would allow factoring out these obfuscations. This is usually taken care of using a “fuzzy” RDD estimator that adjusts the “sharp” effect estimate by incorporating an estimate of the treatment probability $\Pr(D_{i}=1\,|\,G_{i})$, effectively estimating a complier effect.
@@
We, on the other hand, categorize the unit behavior with respect to the assessed rules (G,D)(G,\,D) into complier, defier, nevertaker, alwaystaker and indecisive items. Using this categorization, we derive assumptions for the identification of a complier effect regarding (G,D)(G,\,D). We investigate conditions under which non-change units (e.g., nevertaker), can be disregarded avoiding some of the obfuscations. In the example, units with $X_{2,i}\leq c_{2}$ always have $T_{i}=0$ regardless of $G_{i}$ and thus should have no influence on the complier effect of $G$. In certain cases, this resulting sub-setting principle can justify the usage of “sharp” RDD estimators, which are typically more stable than their “fuzzy” counterparts. Furthermore, we show how the unit categorization changes when progressing along the hierarchy from simple to complex rules incorporating knowledge about the in between categorizations.

### v2

@@ Abstract
RDD (Regression discontinuity design) is a widely used framework for identifying and estimating causal effects at the cutoff of a single running variable. In practice, however, decision-making often involves multiple thresholds and criteria, especially in production systems. Standard MRD (multi-score RDD) methods address this complexity by reducing the problem to a one-dimensional design. This simplification allows existing approaches to be used to identify and estimate causal effects, but it can introduce non-compliance by misclassifying units relative to the original cutoff rules. We develop theoretical tools to detect and reduce “fuzziness” when estimating the cutoff effect for units that comply with individual subrules of a multi-rule system. In particular, we propose a formal definition and categorization of unit behavior types under multi-dimensional cutoff rules, extending standard classifications of compliers, alwaystakers, and nevertakers, and incorporating defiers and indecisive units. We further identify conditions under which cutoff effects for compliers can be estimated in multiple dimensions, and establish when identification remains valid after excluding nevertakers and alwaystakers. In addition, we examine how decomposing complex Boolean cutoff rules (such as AND- and OR-type rules) into simpler components affects the classification of units into behavioral types and improves estimation by making it possible to identify and remove non-compliant units more accurately. We validate our framework using both semi-synthetic simulations calibrated to production data and real-world data from opto-electronic semiconductor manufacturing. The empirical results demonstrate that our approach has practical value in refining production policies and reduces estimation variance. This underscores the usefulness of the MRD framework in manufacturing contexts.
@@ 1 Introduction
In many areas of economics and management, decision-making depends on multiple criteria. Examples include targeted marketing, education policies, and political advertising. The complexity of such settings has motivated the development of formal approaches such as multi-criteria decision analysis, which make trade-offs explicit and offer methods for comparing and ranking alternatives. Once such decision rules are applied in practice, a further challenge is to assess their causal impact. The focus then shifts from choosing among options to evaluating the consequences of those choices. Causal inference, and in particular regression discontinuity design (RDD), provides tools for this purpose.
@@
RDD is a well-established quasi-experimental approach for identifying causal effects whenever treatment assignment changes discontinuously at a threshold in an observed running variable. It has been applied in diverse fields, including economics (Hartmann et al., 2011; Card et al., 2015; Flammer, 2015; Calvo et al., 2019), public policy (Lee, 2008) and the social sciences (Angrist and Lavy, 1999). The appeal of RDD lies in its ability to produce credible causal estimates without relying on the assumptions that are required for most causal inference approaches. In particular, it does not require unconfoundedness Rubin (1974) or global positivity (Austin, 2011), but instead depends on continuity of potential outcomes and the absence of precise manipulation around the cutoff.
@@
The classic RDD exploits a discontinuity in a single treatment assignment rule that depends on an observed running or score variable. When correctly specified, this setting enables average treatment effects to be identified for units near the cutoff. Under mild conditions, it further allows counterfactual reasoning about the decision rule in a small neighborhood around the boundary (Cattaneo et al., 2019; Imbens and Lemieux, 2008; Lee and Lemieux, 2010).
@@
As noted above, treatment decisions in many real-world contexts are based on multiple criteria, particularly in industrial, operational, and engineering systems (Sabaei et al., 2015). The natural extension of the classical RDD is multi-score RDD (MRD). In MRD, treatment depends on a joint rule requiring that multiple score variables satisfy specified threshold conditions simultaneously. The idea was first proposed by Papay et al. (2011). Influential surveys that develop estimators and examine them empirically include Reardon and Robinson (2012), Wong et al. (2013) and Porter et al. (2017). MRD has been widely applied, especially in test-score-based settings (An et al., 2024) and geographical applications (Keele and Titiunik, 2015). However, strategies for identifying MRD effects when multiple cutoffs are combined through arbitrary Boolean rules are still underdeveloped in the literature.
@@
This paper addresses the identification and estimation challenges that arise in multi-score RDD settings. We develop new results tailored to these contexts. In particular, we introduce a formal framework that defines and categorizes behavioral types (“compliers”, “alwaystakers”, “nevertakers”, “defiers”) and analyzes how these types transform when treatment assignment rules change. We investigate general Boolean cutoff mechanisms, such as “AND-type” and “OR-type” rules, which are commonly observed in MRD applications, and analyze how their properties influence local identification. Further, we demonstrate how the complier effect can be identified using these categories. Finally, we illustrate our results with data from a light-emitting diode (LED) production facility, in which each batch is evaluated based on several metrics, such as mean color point measurements or the proportion of units within a lot that meet specified quality requirements.
@@
Our application of the proposed strategies to data from the LED manufacturing industry demonstrates the usefulness of the framework in a real-world setting where treatment assignment depends on multiple quality indicators. We combine the framework with recently developed estimators that use machine learning adjustments to further reduce variance in estimation (Noack et al., 2024). Estimating the causal effect of threshold-based production decisions provides guidance for tuning decision thresholds to improve manufacturing outcomes. In practice, managers need to understand how marginal adjustments to these thresholds affect outcomes such as yield rates and thus overall production efficiency. We complement the empirical analysis with a simulated study based on a semi-synthetic replica of the production environment, highlighting the value of multi-score RDD for counterfactual analysis and manufacturing policy design.
@@
This study contributes to the operations research literature by linking advances in causal machine learning with practical decision-making in manufacturing systems. Our theoretical results on effect identification in MRD create new opportunities for its application in decision-making contexts. By accommodating complex, multi-criteria assignment mechanisms, the framework reflects the realities of operational decision-making practice and provides tools for more informed and effective policy design.
@@
The paper is organized as follows. Section 2 reviews related literature on RDD and its extensions to multi-score settings. Section 3 presents our theoretical framework for identification under general Boolean assignment rules. Section 4 describes the empirical application in opto-electronic semiconductor manufacturing. Section 5 reports results from synthetic and semi-synthetic environments. Section 6 discusses findings from the real-data application, emphasizing the implications for threshold calibration and production efficiency. Section 7 concludes.
@@ 2 Related Literature
RDD dates back to a study on scholarship programs by Thistlethwaite and Campbell (1960). In the following decades, the approach gained popularity in the social sciences, with notable applications by Angrist and Lavy (1999) and Black (1999). Building on this early empirical work, Hahn et al. (2001) provided the first formalized treatment of RDD within the potential outcomes framework and established non-parametric identification and estimation results for both sharp and fuzzy designs, firmly establishing RDD as a core method in modern econometrics.
@@
Important surveys, including those by Imbens and Lemieux (2008), Lee and Lemieux (2010) and Cattaneo et al. (2019), synthesize guidance on estimation choices in RDD, such as bandwidth selection, polynomial order, and diagnostic checks. More recently, several theoretical advances have refined the method. Calonico et al. (2014) developed bias-corrected point estimators and robust confidence intervals that improve coverage accuracy and have become the empirical default for RDD inference. Calonico et al. (2019b) examined the inclusion of covariates in local regression to reduce variance, while Calonico et al. (2019a) analyzed optimal bandwidth selection for robust estimators. Further extensions incorporate machine-learning-based covariate adjustments to further reduce variance, integrating RDD into the broader toolbox of causal machine learning (Kreiss and Rothe, 2022; Arai et al., 2025; Noack et al., 2024).
@@
The idea of extending RDD to multi-score settings dates back to Papay et al. (2011), who showed that causal effects can be identified along a two-dimensional cutoff frontier when program eligibility depends jointly on multiple scores. Reardon and Robinson (2012) systematically collected and formalized MRD identification strategies for cases in which running variables jointly cross their cutoffs to form a multidimensional frontier, corresponding to an “AND”-type rule. They introduced frontier-projection and distance-based estimators, and highlighted the steep data-density and bandwidth challenges that arise as the dimensionality of the assignment space grows. Wong et al. (2013) advanced this line of work with extensive simulation studies, demonstrating that estimators that weight observations along the entire cutoff frontier achieve lower bias and higher efficiency than naïve one-dimension-at-a-time approaches. Their findings underscore the importance of exploiting the joint assignment rule in practice. In another large-scale study, Porter et al. (2017) reached similar conclusions and provided practical recommendations for MRD applications. A related special case, geographic RDD, was introduced by Keele and Titiunik (2015) to evaluate policies along irregular geographic borders.
More recent research has focused on theoretical advances. Choi and Lee (2018) allowed for partial treatment effects when only some running variables cross their cutoffs and introduced thin-plate spline estimators to flexibly recover heterogeneous impacts along a high-dimensional frontier. Building upon earlier work by Imbens and Zajonc (2009), Choi and Lee (2023) analyzed two-dimensional MRD under limited compliance and examined the consequences for estimating complier effects in fuzzy MRD.
@@
Research has also produced more sophisticated MRD estimators. (Imbens and Wager, 2019) proposed a minimax approach, (Liu and Qi, 2024) introduced decision-tree methods, and (Sawada et al., 2025) developed fully multivariate local-polynomial estimators. Cattaneo et al. (2025) examined the statistical properties of important MRD estimators in detail and provided practical recommendations.
@@
RDD has additionally been applied in a variety of business and management contexts. (Hünermund et al., 2021) studied its use in business decision-making, while Ho et al. (2017); Mithas et al. (2022) examined applications in production and operations management. Calvo et al. (2019) used a one-dimensional fuzzy RDD to estimate the effect of procurement officers on budget overruns and delays in public infrastructure projects. Leung et al. (2020) analyzed the effects of labor unionization on the operating performance of supplier firms. In the automobile sector, Hu et al. (2021) employed RDD to assess the impact of regulatory changes on compliance. To the best of our knowledge, however, neither RDD nor MRD has yet been directly applied to production policies.
@@
This paper contributes to the literature in two main ways. First, we derive new theoretical results on unit categorization and effect identification in MRD, which are particularly useful in settings with complex, multi-dimensional decision rules. Secondly, we apply these results to real data, demonstrating the applicability of MRD in operations policy-making and providing evidence to guide actual production policy.
@@ 3 General Identification Strategies in Multi-score RDD
In this section, we first review the basic elements of the RDD and its extension to the multi-score case. We then introduce a formal definition of common unit types (e.g., compliers, defiers) in multi-score, two-stage decision settings that employ cutoff rules for treatment assignment. Using this categorization, we derive an identification result and show that, under certain assumptions, identification does not depend on subsets of unit types with constant response. Finally, we derive rules describing how unit classifications change as assignment rules become more complex.
@@ 3.1 Regression Discontinuity Design
RDD builds on the groundwork of Hahn et al. (2001), which provides a formal justification for effect identification and estimation. An RDD consists of three main elements: a score $X$ that measures a characteristic of the units, a cutoff $c$ that splits the support of the score into two groups, and a treatment $D$ that is assigned to each unit depending on whether its score lies above or below the cutoff, $D_{i}=\operatorname{\mathbbm{1}}[X_{i}\geq c]$ (Cattaneo et al., 2019).
Typically, the parameter of interest is the average treatment effect at the cutoff $c$, $\mathbb{E}[Y(1)-Y(0)\,|\,X=c]$, where $Y(1)$ and $Y(0)$ denote the potential outcomes under treatment and control, respectively Rubin (2005). Identification of this effect relies on the following key assumption:
@@
The conditional mean of the potential outcomes, $\mathbb{E}[Y_{i}(d)\mid X_{i}=x]$ for $d\in\{0,1\}$, is continuous at the cutoff $c$.
This assumption implies that units with scores close to but on opposite sides of the cutoff are comparable in all relevant aspects (Hahn et al., 2001). To rule out selection on gains, it is further required that units near the threshold cannot perfectly manipulate their scores. Formally, for sufficiently small $\epsilon>0$, $\mathbb{E}[Y(1)-Y(0)\mid D,\,X=x]=\mathbb{E}[Y(1)-Y(0)\mid X=x]$ must hold for $x\in(c-\varepsilon,c+\varepsilon)$ (Chernozhukov et al., 2024).
Under these relatively mild assumptions, RDD provides inference around the threshold that is as credible as inference from a randomized experiment (Lee, 2008). The average treatment effect at the cutoff,
@@ 3.2 Problem description
While standard RDD is suitable for analyzing the effect of a single score variable at a cutoff, real-world treatment decisions often depend on multiple score variables. For example, treatment $T_{i}=1$ may be assigned to unit $i$ only if both criteria $X_{1,i}>c_{1}$ and $X_{2,i}>c_{2}$ are met. Such multi-score cases are often reduced to a standard RDD by transforming the variables into a single composite score, allowing identification and estimation to be justified by results from the one-dimensional case. A drawback of this approach, however, is the reduced interpretability of the transformed score. This raises the question of the precise conditions under which a treatment effects exist and can be identified in a multi-score setting. In particular, a well-defined multidimensional cutoff effect requires some independence of the directions from which the cutoff is approached.
@@
In applications, the aim is usually to optimize the treatment assignment $T_{i}$ in order to improve the outcome $Y_{i}$. In some critical systems, however, only gradual adjustments to $T_{i}$ are feasible. In such cases, part of the assignment rule $T_{i}$ may be treated as fixed; for example $H_{i}\coloneqq\operatorname{\mathbbm{1}}[X_{2,i}>c_{2}]$ is accepted as given, while the remaining part, that is $G_{i}\coloneqq\operatorname{\mathbbm{1}}[X_{1,i}>c_{1}]$, is subject to optimization. The effect of a subrule $G$ at its cutoff can therefore provide insights into how the overall assignment rule $T$ might be improved. In some settings, the assignment rule $T$ serves only as a recommendation, and the actual treatment $D$ may deviate from it. In the RDD literature, this case is referred to as “fuzzy”. Altogether, this leads to a hierarchy of decision rules $G\to T\to D$. Each step in the hierarchy obscures the assessment of $G$ (for example, evaluation of the optimal value of cutoff $c_{1}$) because at every level of it a unit may become non-compliant. Knowledge of the unit’s compliance type with respect to (G,D)(G,\,D), however allows these ambiguities to be addressed. In practice, this is usually handled using a fuzzy RDD estimator, which adjusts the sharp effect estimate by incorporating an estimate of the treatment probability $\Pr(D_{i}=1\,|\,G_{i})$, thereby recovering a complier effect.
@@
In contrast, we categorize unit behavior with respect to the rules (G,D)(G,\,D) into the standard compliance types: compliers, defiers, nevertakers, alwaystakers, and an additional category of indecisive units. Based on this categorization, we derive assumptions for identifying the complier effect with respect to (G,D)(G,\,D). We then examine conditions under which non-change units (e.g., nevertakers) can be excluded without affecting identification and thus avoiding some of the ambiguities. For example, units with $X_{2,i}\leq c_{2}$ always have $T_{i}=0$ regardless of $G_{i}$ and thus may not contribute to the complier effect of $G$. In some cases, this subsetting principle justifies the use of sharp RDD estimators, which are typically more stable than their fuzzy counterparts. Finally, we show how the unit classifications change when moving on the hierarchy from simple to more complex assignment rules, incorporating knowledge of intermediate categorizations.